OpenStudy (anonymous):

Find the equation of a parabola with focus F(0, -3) and directrix y = 3.

OpenStudy (anonymous):

@mml

OpenStudy (anonymous):

haha you love tagging me dont you ?

OpenStudy (tkhunny):

Do you know what it looks like? (y-k) = p(x-h)^2 or (x-h) = p(y-k)^2 You must know this or you can't do it very well.

OpenStudy (anonymous):

Yes! :P

OpenStudy (tkhunny):

Alternatively, the Definition of the Parabola can be useful. Locus of points equidistant from a given point (focus) and a given line (directrix). Given all points, (x,y) The distance from (0,-3) is \(\sqrt{(x-0)^{2} + (y+3)^{3}} = \sqrt{x^{2} + y^{2} + 6x + 9}\) The distance from y = 3 is \(|y-3|\) Those two expressions must be equal. \(\sqrt{x^{2} + y^{2} + 6x + 9} = |y-3|\) \(x^{2} + y^{2} + 6x + 9 = y^{2} - 6y + 9\) You're almost done.